Since most people will probably not attempt to read the RHIC e-logs I forward the todays summary indicating what might happen during the night and morning shifts Flemming ------------------------------------------------------ Flemming Videbaek Physics Department Brookhaven National Laboratory tlf: 631-344-4106 fax 631-344-1334 e-mail: videbaek@bnl.gov The gammaT work is done, and operations crews are starting to sweep RHIC. With limited crew and many important support activities ongoing throughout the day, a lot of sweeping remains to be done. The blue ring is tuned on upramp to slow factor 5, and this tuning will probably hold to slow factor 2, though it crashes on downramp. Yellow ramp tuning has been compromised by a failure of the yo8-qd3 power supply -- Don Bruno is replacing the backplane for this unit. Tuning in both rings cannot continue until this repair is done. It is unlikely that we will have 100 GeV/u ramps tonight; if not, George plans to leave us 65 GeV/u ramps for overnight. There have been many small things but no unaticipated showstoppers in the 100 GeV/u tuneup, which will continue into Sun Jul 15. ETA for return to beam tonight remains sometime late evening, perhaps 2200-2300. Carl Schultheiss and controls are working around the 720 Hz glitch problem on the main power supplies; it is confirmed that they are associated directly with the ramp spline control points, as hypothesized by Al Marusic earlier today. --------------------------------------------------------------------------------
This archive was generated by hypermail 2b30 : Sat Jul 14 2001 - 20:23:12 EDT